OpenStudy (anonymous):

find the exact value of cos 20pi/3

OpenStudy (anonymous):

i tried this problem and cant figure it out someone help please?

OpenStudy (waleed_imtiaz):

yea its simple ...cos 20pi/3 = cos1200 =cos [14*90 -60]=-cos60=-1/2

OpenStudy (anonymous):

cos20pi/3=cos2pi/3=-1/2